OpenStudy (mendicant_bias):

Alright, another Dirac-Delta. Getting the hang of this. Posted below in a moment.

OpenStudy (mendicant_bias):

\[y''-y=\delta(t-2\pi),\ \ \ y(0)=0, \ y'(0)=1\]

OpenStudy (mendicant_bias):

\[[s^{2}Y(s)-sy(0)-y'(0)]-Y(s)=e^{-2\pi s}\]

OpenStudy (mendicant_bias):

\[(s^2-1)Y(s)=e^{-2\pi s}+1\]

OpenStudy (mendicant_bias):

\[Y(s)=\frac{e^{-2\pi s}+1}{s^2-1}\]

OpenStudy (mendicant_bias):

Going to separate the numerator and evaluate them separately

OpenStudy (sidsiddhartha):

yup :)

OpenStudy (mendicant_bias):

\[\frac{e^{-2\pi s}}{s^2-1}+\frac{1}{s^2-1}=\sinh(t-2\pi)\mathcal{U}(t-2\pi)+\sinh(t)\]

OpenStudy (mendicant_bias):

Looking over that, not sure about it

OpenStudy (sidsiddhartha):

correct :) well done again

OpenStudy (mendicant_bias):

Eghhh, I think we made a mistake somewhere, heh, yeah, we (I) did

OpenStudy (sidsiddhartha):

where?

OpenStudy (mendicant_bias):

Answer is close, but not identical\[y=\sin(t)+\sin(t)\mathcal{U}(t-2\pi)\]

OpenStudy (mendicant_bias):

I'm not sure

OpenStudy (sidsiddhartha):

another typo in your book they missed that hyperbolic again

OpenStudy (mendicant_bias):

I made a mistake, too! It was my fault, it was y'' + y

OpenStudy (mendicant_bias):

Good thing is, I understand the process and am getting it, w/o typos I am doing things right, just need to make sure I cover all of my bases

OpenStudy (sidsiddhartha):

oh ok now we are talking :P

OpenStudy (mendicant_bias):

Heh, alright, taking a look at it again before I leave it, shouldn't take too long to fix

OpenStudy (mendicant_bias):

What I still don't understand (I can understand the change from them to sine trig functions, is why the sine function attached to the unit step function term doesn't have its argument altered in the problem.

OpenStudy (sidsiddhartha):

yes argument of the sine have to be altered by 2pi

OpenStudy (mendicant_bias):

Alright, col! Going to move on to another problem.

OpenStudy (sidsiddhartha):

okaay :)